South Shore Construction builds permanent docks and seawalls along the southern shore of Long Island, New York. Although theр үү 2 3 4 5 6 7 8 9 | 1 L-501 10 11 12 13 14 15 16 17 18 19 20 Time Period (t) - Select your answer - What type of patternPlease help I am very confused on how to solve this
Answer #1

Solve in Excel, input data in the following manner:

Year Quarter Period Q1 Q2 Q3 Q4 Sales
1 1 1 1 0 0 0 21
1 2 2 0 1 0 0 97
1 3 3 0 0 1 0 172
1 4 4 0 0 0 0 22
2 1 5 1 0 0 0 10
2 2 6 0 1 0 0 111
2 3 7 0 0 1 0 215
2 4 8 0 0 0 0 23
3 1 9 1 0 0 0 60
3 2 10 0 1 0 0 154
3 3 11 0 0 1 0 329
3 4 12 0 0 0 0 48
4 1 13 1 0 0 0 92
4 2 14 0 1 0 0 197
4 3 15 0 0 1 0 389
4 4 16 0 0 0 0 83
5 1 17 1 0 0 0 163
5 2 18 0 1 0 0 292
5 3 19 0 0 1 0 439
5 4 20 0 0 0 0 181

Sales 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 Time Period

a) Horizontal trend

b) Using Excel, go to Data, select Data Analyis, choose Regression. Put Q1, Q2, Q3 and Q4 in X input range and Sales in Y input range.

Coefficients Standard Error t Stat P-value
Intercept 71.4 36.93891834 1.932921 0.072351
Q1 -2.2 52.2395193 -0.04211 0.966964
Q2 98.8 52.2395193 1.891288 0.078057
Q3 237.4 52.2395193 4.544452 0.000387
Q4 0 0 65535 #NUM!

Revenue = 71 - 2*Q1 + 99*Q2 + 237*Q3

Using above equation:

Q1 Forecast = 71 - 2*1 + 99*0 + 237*0 = 69

Q2 Forecast = 71 - 2*0 + 99*1 + 237*0 = 170

Q3 Forecast = 71 - 2*0 + 99*0 + 237*1 = 308

Q4 Forecast = 71 - 2*0 + 99*0 + 237*0 = 71

c) Using Excel, go to Data, select Data Analyis, choose Regression. Put Q1, Q2, Q3, Q4 and Period in X input range and Sales in Y input range.

Coefficients Standard Error t Stat P-value
Intercept -73.2 21.0610256 -3.47561 0.003711
Period 12.05 1.277012252 9.436088 1.91E-07
Q1 33.95 20.78825335 1.633134 0.124721
Q2 122.9 20.59120385 5.968568 3.44E-05
Q3 249.45 20.47206377 12.1849 7.68E-09
Q4 0 0 65535 #NUM!

Revenue = -73.2 + 12.05*Period + 33.95*Q1 + 122.9*Q2 + 294.45*Q3

Using above equation:

Q1 Forecast = -73.2 + 12.05*21 + 33.95*1 + 122.9*0 + 294.45*0 = 213.8

Q2 Forecast = -73.2 + 12.05*22 + 33.95*0 + 122.9*1 + 294.45*0 = 314.8

Q3 Forecast = -73.2 + 12.05*23 + 33.95*0 + 122.9*0 + 294.45*1 = 453.4

Q4 Forecast = -73.2 + 12.05*24 + 33.95*0 + 122.9*0 + 294.45*0 = 216